A 4% annual coupon convertible bond has 14 years to maturity, a $1,000 par value and is priced at $1,620. This bond can be converted into 20 shares of common. The common currently sells for $63 per share and pays a $0.20 quarterly dividend per share. Similar bonds that are not callable yield 5.25%.

A) What is the bonds conversion equivalent?

B) What is the bond's payback period?

C) What is this bond's investment value?
Answer #1

1.
=20*63=1260

2.
We see that the bond's payback period is given as equal to=13+(1620-4%*1000*13)/(1000+4%*1000)=14.0577

3.
=Coupon rate*Par value/yield*(1-1/(1+yield)^t)+Par value/(1+yield)^t
=4%*1000/5.25%*(1-1/1.0525^14)+1000/1.0525^14
=878.2212